# Perturbation-based Effect Measures for Compositional Data

### v4

@@ 3.1 Efficient estimation using semiparametric theory
We now give a brief introduction to semiparametric estimation of statistical functionals and, in particular, we describe how naive plug-in estimation of target parameters based on nonparametric methods rarely leads to efficient confidence intervals and how to remedy this. This perspective is not new and goes back to at least Pfanzagl (1982) with several more recent monographs and reviews (Bickel et al., 1993; Tsiatis, 2006; Chernozhukov et al., 2018; Kennedy, 2023). The recent surge in the popularity of these methods is fueled by the advent of powerful machine learning methods that use flexible black box models to learn the nuisance functions involved in the estimation of functionals of interest. For illustration, we focus on the estimation of $\tau_{\psi}$, that is, the estimation of
$\tau_{\psi}=\mathbb{E}[\partial_{\ell}f(\ell,W)\bigm{|}_{\ell=L}],$
where $f:(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$. A naive approach, which is known as plug-in estimation, is to estimate $\tau_{\psi}$ by first learning a differentiable estimate $\widehat{f}$ of $f$ using all $n$ observations and then computing
$\widetilde{\tau}_{\psi}:=\frac{1}{n}\sum_{i=1}^{n}\partial_{\ell}\widehat{f}(L_{i},W_{i}).$
While such an estimate can be consistent, it is rarely asymptotically Gaussian, which invalidates standard asymptotic inference. Using the corresponding plug-in estimate of the asymptotic variance, given by
@@
3.2 Efficient estimation of perturbation effects
In this section, we present our proposals for estimating perturbation effects. Further details and full algorithms can be found in Section S4.1. All the presented estimators involve regressions that can be performed by arbitrary (sufficiently predictive) regression methods.
3.2.1 Nonparametric estimation of the binary perturbation effect
The denominator and $\mathbb{E}[Y]$ are both easy to estimate efficiently using their empirical counterparts (the one-step correction term is 00 for these quantities). The term $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ in the numerator is trickier and is related to the notion of the average predictive effect (e.g., Robins et al., 1994; Chernozhukov et al., 2018; Kennedy, 2023) that has been analyzed extensively in the causal inference literature. The one-step corrected estimator for $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ corresponds to the augmented inverse propensity weighted (AIPW) estimator and is given by
$\widehat{\lambda}_{\psi}:=\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{f}(1,W_{i})-Y_{i}+\frac{Y_{i}-\widehat{f}(L_{i},W_{i})}{\widehat{\pi}(W_{i})}L_{i}}{\frac{1}{n}\sum_{i=1}^{n}(1-L_{i})}.$ (10)
3.2.2 Nonparametric estimation of the directional perturbation effect
This quantity is another well-known target of inference; the average partial effect (Newey and Stoker, 1993; Chernozhukov et al., 2022; Klyne and Shah, 2023). The one-step corrected estimator in this setting is given by
$\widehat{\tau}_{\psi}:=\frac{1}{n}\sum_{i=1}^{n}\partial_{\ell}\widehat{f}(L_{i},W_{i})-\widehat{\rho}(L_{i},W_{i})(Y_{i}-\widehat{f}(L_{i},W_{i}))$ (11)
For the estimation of $f$ we require that the estimator $\widehat{f}$ is differentiable and with computable derivatives (which additionally need to approximate the true derivative at a sufficiently fast rate). This excludes the direct use of many popular machine learning methods such as random forests or boosted trees. One solution is to start from a potentially non-differentiable pilot estimate of $f$ and smooth it to get a differentiable estimate for which derivatives can be computed. Such a procedure based on kernel-smoothing has been proposed by Klyne and Shah (2023). Here, we use an alternative smoothing method which uses random forests and local polynomials. Finally, for the estimation of $\rho$, we follow the proposal in Klyne and Shah (2023) and assume a location-scale model. This requires the estimation of the conditional mean and variance of $L$ given $W$ (referred to as $\eta$ in Algorithm S1). Further details on our smoothing method and the location-scale based estimate of $\rho$ are given in Sections S4.5 and S4.6, respectively.
3.2.3 Partially linear estimation of perturbation effects
While the estimators in the previous sections make only mild assumption on the data-generating model, they rely on accurate estimates of quantities that may be difficult to estimate, such as derivatives of densities and regression functions (see Section S4.3 for an illustration of this sensitivity). One way to avoid these complicated estimations is by additionally assuming a partially linear model. For both binary and directional perturbations (with a derivative-isolating reparametrization), we can assume a partially linear model on $L$ and $W$, that is, the existence of $h$ and $\theta$ such that
$\mathbb{E}[Y\,|\,L,W]=\theta L+h(W).$ (12)
With this assumption, $\theta$ is exactly equal to the perturbation effect. Efficient estimation of $\theta$ in a partially linear model is a well-studied problem (e.g., Robinson, 1988; Chernozhukov et al., 2018) and one estimator (sometimes referred to as the partialling out estimator) is given by
@@ 3.2.3 Partially linear estimation of perturbation effects
where $\widehat{g}$ is an estimator of $g:w\mapsto\mathbb{E}[Y\,|\,W=w]$ and $\widehat{m}$ is an estimator of $m:w\mapsto\mathbb{E}[L\,|\,W=w]$. In contrast to the estimators from the previous sections, this estimator only relies on estimating two conditional mean functions. This comes at the expense of assuming that there is no interaction effect between $L$ and $W$ on $Y$ and that the effect of $L$ is linear on $Y$. Even when these assumptions are not satisfied, the estimator may still provide a reasonable estimate of the effect of $L$ on $W$ and we found in our experiments that the estimator is more robust to errors in the estimation of the nuisance functions than the estimators above. In fact, one can show that the quantity $\mathbb{E}[(\mathbb{E}[Y\,|\,L,W]-\theta L-h(W))^{2}]$ is minimized by the population version of the partially linear model estimate, so one can think of the partially linear model as a ‘best approximation’ to a generic nonparametric model (see Proposition S4 in Section S4.3.3 for a proof).
@@ 4.1.1 Effects of diversity on income in semi-synthetic US census data
Based on the two constructed community-level datasets, we now compare a naive approach that neglects the simplex structure and is applied in the literature (e.g. Richard (2000); Richard et al. (2007); Lu et al. (2015); Paarlberg et al. (2018)) with the perturbation-based approach. The naive approach summarizes racial diversity in a single measure, here we use one minus the Gini coefficient $L=1-\frac{1}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|Z^{j}-Z^{k}|$ as in Example 2, for which large values correspond to high diversity, and estimate associations between this measure and the average compensation $Y$. We call the method that regresses $Y$ on $L$ linearly naive_diversity. This estimator can easily be adapted to also adjust for the additional covariates $X$, the racial composition $Z$ or both under a partially linear model assumption. The point estimates together with asymptotic confidence intervals (based on an efficient partialling out estimator like the estimator in Section 3.2.3) are provided in Table 2 where random forests are used for nuisance function estimation when including additional covariates. The naive estimators that condition either on the empty set or on $X$ both result in significant negative effects which by construction of the datasets is incorrect. Furthermore, additionally conditioning on $Z$, leads to an ill-conditioned estimator as $Z$ is now a perfect predictor of the diversity measure. This is also visible in the large resulting confidence intervals. While we here constructed the groups in a way to highlight how compositional confounding can be misleading, we believe that this type of confounding is likely also present in real-world settings as groupings are often confounded e.g., educational levels may well be correlated with race and average compensation in actual communities.
@@ 4.2 Variable influence measures when predicting BMI from gut microbiome
We begin by evaluating, how sensitive the proposed perturbation effect estimates are with respect to the employed nuisance function estimators. To this end, we consider three different regression methods: random forest (rf), neural network (mlp) and kernelized support vector regression using an RBF kernel (svr). Each of the regressions is combined with a grid-search cross-validation to automatically tune the hyperparameters of each method. For each of the three regression methods and all the species we then compute $\mathrm{CFI}_{\mathrm{mult}}$, $\mathrm{CFI}_{\mathrm{unit}}$ and $\mathrm{CKE}$ using the estimator based on a partially linear model and the $\mathrm{CKE}$ using the nonparametric estimator. As baselines, we additionally compute a nonparametric $R^{2}$ (Williamson et al., 2021) and permutation importance (Breiman, 2001). Figure 2 shows the Spearman correlation between the estimates from the different regression methods.
Figure 2: Spearman correlations of variable importance measures computed on each feature using different regressions. On the left, these are computed on all $561$ features and on the right they are computed on the subset of features where at least one of the two methods deems the effect significant at a $5\%$ level. The $R^{2}$ and permutation methods do not come with significance tests and are therefore omitted from the right plot.
While the average perturbation effect estimates which are specifically tailored to the compositional setting are consistent across different regression methods, the permutation and $R^{2}$ measures are – unsurprisingly – not. The population version of the nonparametric $R^{2}$ is always 00, so the estimates are essentially mean zero random noise that depends on the estimator. For the permutation measure the inconsistency comes from the fact that the permuted data inherently leave the support of the simplex, making the resulting influence measure highly dependent on how the regression methods extrapolate. Unfortunately, due to a lack of alternatives, it is still common to use these types feature importance or influence measures in microbiome science (e.g., Marcos-Zambrano et al., 2021). Furthermore, the results show that the nonparametric estimator of the CKE is less well-behaved than the other estimators based on the partially linear model. One reason for this could be that dividing by the propensity score estimator results in a more unstable estimate, particularly when the propensity score is poorly estimated.
@@
We now analyze different types of average perturbation effects and compare them with the commonly used log-contrast based analysis. We fit both a standard log-contrast regression and an $\ell^{1}$-penalized log-contrast regression and then use the coefficients as feature influences. Since the log-contrast model is only defined on the open simplex, that is, it does not permit zeros in the data, one usually adds a ‘pseudocount’ to each observation before rescaling to the simplex. We follow a standard recommendation and use the minimum non-zero observation of $Z$ divided by $2$ (this is not required for the perturbation-based measures). To illustrate the difference between the approaches, we present the $p$-values from $\mathrm{CFI}_{\mathrm{mult}}$, $\mathrm{CFI}_{\mathrm{unit}}$, $\mathrm{CKE}$ and the log-contrast method for several species in the left of Figure 3. From these $p$-values, we can see that the log-contrast and $\mathrm{CKE}$ are highly correlated and appear to be selecting similar species. This is despite the fact that the log-contrast model is supposed to model the continuous effect of each species. The similarities to the $\mathrm{CKE}$ hence indicates (likely due to the pseudocounts) that the log-contrast coefficients are instead capturing the effects of presence and absence of species. To further investigate this, for different values of pseudocounts we plot log-contrast estimates (scaled by the standard deviation of $\log(Z^{j}+\text{pseudocount})$) in the right of Figure 3. This further highlights the dependence of the log-contrast coefficients on the used pseudocount. The fact that the perturbation-based effects explicitly account for zeros and allow us to distinguish between continuous effects (e.g., $\mathrm{CFI}_{\mathrm{unit}}$ or $\mathrm{CFI}_{\mathrm{mult}}$) and presence/absence effects $\mathrm{CKE}$ are an important feature of the perturbation-based approach. In Section S5.7 we provide additional plots which compare the significant hits of $\mathrm{CFI}_{\mathrm{mult}}$, log-contrast estimates and a naive marginal analysis and illustrate the differences between $\mathrm{CFI}_{\mathrm{unit}}$ and $\mathrm{CKE}$ when compared to estimates using $\mathrm{CFI}_{\mathrm{mult}}$.
@@ Acknowledgements
Table S1: Numerical illustration of Examples S1 and S2. The first row in each table corresponds to the marginal treatment effect (MTE or OLS of $Y$ on $-G(Z)$) estimate based on a correctly specified linear model. The second row correspond to estimates of two of the target parameters introduced in Section 2.2 with random forest (rf) as nuisance function estimators. Bold red numbers indicate a significant negative effect at a $5\%$ level, while bold green indicates a significant positive effect. Left: The estimated MTE is negative, even though the true effect is positive. The compositional knock-out effect ($\mathrm{CKE}^{1}$), which adjust for the simplex structure, results in the correct sign. Right: The estimated marginal effect of an increase in diversity (measured by the negative Gini coefficient, i.e., $-G(Z)$) is negative even though the true effect is equal to $1$ by construction. The compositional diversity effect $\mathrm{CDI}_{\mathrm{Gini}}$ correctly recovers the effect. Details on both experiments are provided in Section S5.1 of the supplementary material.
@@
In this section we provide a general method for finding derivative-isolating reparametrizations based on the development in Section 2.1.2. Our main result is as follows.
@@
Part (a) starts from a given speed and finds a corresponding derivative-isolating reparametrization by solving the differential equation in (S3). Part (b) starts from a one-dimensional summary statistic, which is assumed to be smoothly decreasing in the distance from $z$ to the endpoint. Using the statistic as one component in a derivative-isolating reparametrization, the result provides the speed of the perturbation that is being parametrized. In the following two examples, we give the missing details from Examples 1 and 2 from Section 2.1.2.
@@
S2.6 Amalgamations
In this section we consider general amalgamating perturbations as defined in Section 2.1.1 rather than only those which push towards a vertex of the simplex, i.e., those we describe in Section 2.2.1. We first generalize Example 1.
@@
With this generalization, we can consider perturbation effects for general amalgmating perturbations. As an example, consider a study where data is collected on how individuals spend their time each day (e.g., hours spent sleeping, eating, working, doing sports, etc.) as well as a response variable (e.g., general happiness). Assume now that we want to use this data to investigate the relationship between time spent on sports and happiness. As in Section 2.2.1, we are again interested in the effect of an individual component, so we might consider estimating a $\mathrm{CFI}^{j}$. However, this would capture the effect of increasing time spent on sports while all other activities are down-scaled proportionally. For example, time spent sleeping and eating would both be decreased if time doing sports is increased which is not realistic. To address this, we can instead consider a general $(A\to B)$-amalgamating perturbation (as defined in Section 2.1.1), where $B=\{\text{sport}\}$ and $A$ corresponds to activities which are reduced to make increased time for sports. Perturbing in this fashion keeps the components in $[d]\setminus(A\cup B)$ fixed (e.g., eating and drinking) while down-scaling the activities in $A$. We call directional perturbation effects of this form compositional amalgamation influences, or $\mathrm{CAI}^{A\to B}$ for short. As with the $\mathrm{CFI}^{j}$, we use subscripts to denote the speed of the perturbation, so we have $\mathrm{CAI}_{\mathrm{unit}}^{A\to B}$ and $\mathrm{CAI}^{A\to B}_{\mathrm{mult}}$ (see Example S5).
Analogously, we could be interested in the effect of removing particular activities (e.g. gambling) entirely and spending the time on desirable activities (e.g. sports) while fixing time spent on a third group of activities (e.g. eating, sleeping). We can capture this effect using a binary amalgamating perturbation $\psi$. We call the resulting average binary perturbation effect for $\psi$ an $(A\to B)$-compositional amalgamation effect, or $\mathrm{CAE}^{A\to B}$ for short. We summarize these target parameters in Table S2 which complements Table 1.
$\mathrm{CAE}^{A\to B}$ binary $\mathsf{A}_{A\to B}(z)$ $\mathds{1}_{\{z^{A}=0\}}$
Table S2: List of amalgamating perturbation effects considered in this work and what effects they capture. For each perturbation effect the corresponding speed, endpoint and $L$-term of the reparametrization used for efficient estimation is included. All reparametrizations use the canonical direction $v_{E}$ based on the endpoint, see (6). The $W$-term $\phi^{W}(z)$ of the reparametrization is $E(z)$ for the binary effects and $(v_{E}(z),E(z))$ for the directional effects.
@@
In this section we describe an alternative way to construct perturbations than the one considered in Sections 2.1.1 and 2.1.2. When constructing the directional perturbations in Section 2.1.2, we considered (Euclidean) straight lines to the endpoint to construct directions, that is, $v_{E}:z\mapsto(E(z)-z)/\|E(z)-z\|_{1}$. While we believe that this is natural in many applications, one could also imagine starting from alternative geometries on the simplex and construct other types of directions. We illustrate how this would work for the Aitchison geometry and show that it provides an interesting connection to the log-ratio based literature. The Aitchison geometry is a popular alternative geometry that goes back to the work of Aitchison [1982] who proposed to analyze the interior of the simplex by transforming to a subset of Euclidean space and applying conventional methods to the transformed data. A popular variant of this transformation is the centered log-ratio transform (CLR) defined for all $z\in(\Delta^{d-1})^{\circ}$ (the interior of the simplex) by
@@
As a first example, assume that the SCM given in Figure S2 (left) generates the data. Given a perturbation $\psi$ that we are interested in, a derivative-isolating reparametrization (see Section 2.1.2) $\phi:\Delta^{d-1}\rightarrow\mathbb{R}\times\mathcal{W}$ maps $Z$ to $(L,W)$ such that if we intervene on $Z$ by setting it to $\psi(Z,\gamma)$ this corresponds to an intervention in an SCM over the variables $(Y,L,W)$ on $L$ alone (see Figure S2 (right)).
@@
In this section we provide details on the theory and algorithms for the estimation of perturbation effects. Section S4.1 contains a summary of the theoretical results and algorithms for the main estimation procedures. Section S4.2 contains a proof of our main theoretical result. Section S4.3 investigates the robustness of the different estimators. Section S4.4 describes how to modify the estimation of $\tau_{\psi}$ using a derivative-isolating reparametrization in the presence of observations with $\omega_{\psi}(Z)=0$. Section S4.5 describes our procedure for the estimation of the derivative of a generic regression function. Section S4.6 describes our procedure for the estimation of the conditional score function $\rho$ that is required for the estimation of an average partial effect.
S4.1 Theoretical analysis of estimators
In this section we present a result on asymptotic normality uniformly over large classes of distributions for each of the three estimation methods described in Section 3; nonparametric estimation of $\tau_{\psi}$, nonparametric estimation of $\lambda_{\psi}$ and by means of estimating $\theta$ in the partially linear model (12). These results are derived using standard proof techniques from the semiparametric literature and likely exist individually in similar form elsewhere [e.g., Klyne and Shah, 2023, Chernozhukov et al., 2018]. We do not claim that any of the derived asymptotic results are novel or surprising but include them for completeness and to motivate the precise form of the estimators and their corresponding variance estimates.

### v5

@@ 4.1 Efficient estimation using semiparametric theory
We now give a brief introduction to semiparametric estimation of statistical functionals and, in particular, we describe how naive plug-in estimation of target parameters based on nonparametric methods rarely leads to efficient confidence intervals and how to remedy this. This perspective is not new and goes back to at least Pfanzagl (1982) with several more recent monographs and reviews (Bickel et al., 1993; Tsiatis, 2006; Chernozhukov et al., 2018; Kennedy, 2023). The recent surge in the popularity of these methods is fueled by the advent of powerful machine learning methods that use flexible black box models to learn the nuisance functions involved in the estimation of functionals of interest. For illustration, we focus on the estimation of $\tau_{\psi}$, that is, the estimation of $\tau_{\psi}=\mathbb{E}[\partial_{\ell}f(\ell,W)\bigm{|}_{\ell=L}],$ where $f:(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$. A naive approach, which is known as plug-in estimation, is to estimate $\tau_{\psi}$ by first learning a differentiable estimate $\widehat{f}$ of $f$ using all $n$ observations and then computing $\widetilde{\tau}_{\psi}:=\frac{1}{n}\sum_{i=1}^{n}\partial_{\ell}\widehat{f}(L_{i},W_{i}).$ While such an estimate can be consistent, it is rarely asymptotically Gaussian, which invalidates standard asymptotic inference. Using the corresponding plug-in estimate of the asymptotic variance, given by
@@
4.2 Partially linear estimation of perturbation effects
In this section, we present our proposals for estimating perturbation effects. It is possible to estimate average perturbation effects under mild smoothness assumptions entirely nonparametrically. We describe these estimators in Section S4.1. However, they rely on accurate estimates of quantities that may be difficult to estimate, such as derivatives of densities and regression functions (see Section S4.4 for an illustration of this sensitivity). One way to avoid these complicated estimations is by additionally assuming a partially linear model. For both binary and directional perturbations (with a derivative-isolating reparametrization), we can assume a partially linear model on $L$ and $W$, that is, the existence of $h$ and $\theta$ such that $\mathbb{E}[Y\,|\,L,W]=\theta L+h(W).$ With this assumption, $\theta$ is exactly equal to the perturbation effect. Efficient estimation of $\theta$ in a partially linear model is a well-studied problem (e.g., Robinson, 1988; Chernozhukov et al., 2018) and one estimator (sometimes referred to as the partialling out estimator) is given by
@@ 4.2 Partially linear estimation of perturbation effects
where $\widehat{g}$ is an estimator of $g:w\mapsto\mathbb{E}[Y\,|\,W=w]$ and $\widehat{m}$ is an estimator of $m:w\mapsto\mathbb{E}[L\,|\,W=w]$. In contrast to the estimators from the previous sections, this estimator only relies on estimating two conditional mean functions. This comes at the expense of assuming that there is no interaction effect between $L$ and $W$ on $Y$ and that the effect of $L$ is linear on $Y$. Even when these assumptions are not satisfied, the estimator may still provide a reasonable estimate of the effect of $L$ on $W$ and we found in our experiments that the estimator is more robust to errors in the estimation of the nuisance functions than the estimators above. In fact, one can show that the quantity $\mathbb{E}[(\mathbb{E}[Y\,|\,L,W]-\theta L-h(W))^{2}]$ is minimized by the population version of the partially linear model estimate, so one can think of the partially linear model as a ‘best approximation’ to a generic nonparametric model (see Proposition S4 in Section S4.4.3 for a proof). Further details on both the partialling out and nonparametric estimators (including full algorithms) can be found in Section S4.2.
@@ 5.1.1 Effects of diversity on income in semi-synthetic US census data
Based on the two constructed community-level datasets, we now compare a naive approach that neglects the simplex structure and is applied in the literature (e.g. Richard (2000); Richard et al. (2007); Lu et al. (2015); Paarlberg et al. (2018)) with the perturbation-based approach. The naive approach summarizes racial diversity in a single measure, here we use one minus the Gini coefficient $L=1-\frac{1}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|Z^{j}-Z^{k}|$ as in Example 2, for which large values correspond to high diversity, and estimate associations between this measure and the average compensation $Y$. We call the method that regresses $Y$ on $L$ linearly naive_diversity. This estimator can easily be adapted to also adjust for the additional covariates $X$, the racial composition $Z$ or both under a partially linear model assumption. The point estimates together with asymptotic confidence intervals (based on an efficient partialling out estimator like the estimator in Section 4.2) are provided in Table 2 where random forests are used for nuisance function estimation when including additional covariates. The naive estimators that condition either on the empty set or on $X$ both result in significant negative effects which by construction of the datasets is incorrect. Furthermore, additionally conditioning on $Z$, leads to an ill-conditioned estimator as $Z$ is now a perfect predictor of the diversity measure. This is also visible in the large resulting confidence intervals. While we here constructed the groups in a way to highlight how compositional confounding can be misleading, we believe that this type of confounding is likely also present in real-world settings as groupings are often confounded e.g., educational levels may well be correlated with race and average compensation in actual communities.
@@ 5.2 Variable influence measures when predicting BMI from gut microbiome
In Section S5.7 we show that non-compositional variable importance measures such as permutation-based feature importance or nonparametric $R^{2}$ exhibit undesirable properties when applied to this data. In particular, there is little correlation between important features identified by applying different nonparametric regression methods.
As a consequence of these findings, we focus on analyzing different types of average perturbation effects and comparing them with commonly used compositional feature importance measures, namely those based on log-contrast models. The linear log-contrast model assumes that there exists $\beta\in\mathbb{R}^{d}$ such that
$\mathbb{E}[Y\,|\,Z]=\sum_{j=1}^{d}\beta^{j}\log(Z^{j})\quad\text{with}\quad\sum_{j=1}^{d}\beta^{j}=0.$
@@
We fit this model and an $\ell^{1}$-penalized version of the same model and seek to use the coefficients as feature influences. Since the log-contrast model is only defined on the open simplex, that is, it does not permit zeros in the data, one usually adds a ‘pseudocount’ to each observation before rescaling to the simplex. We follow a standard recommendation and use the minimum non-zero observation of $Z$ divided by $2$ (this is not required for the perturbation-based measures). To illustrate the difference between the approaches, we present the $p$-values from $\mathrm{CFI}_{\mathrm{mult}}$, $\mathrm{CFI}_{\mathrm{unit}}$, $\mathrm{CKE}$ and the log-contrast method for several species in the left of Figure 2. From these $p$-values, we can see that the log-contrast and $\mathrm{CKE}$ are highly correlated and appear to be selecting similar species. This is despite the fact that the log-contrast model is supposed to model the continuous effect of each species. The similarities to the $\mathrm{CKE}$ hence indicates (likely due to the pseudocounts) that the log-contrast coefficients are instead capturing the effects of presence and absence of species. To further investigate this, for different values of pseudocounts we plot log-contrast estimates (scaled by the standard deviation of $\log(Z^{j}+\text{pseudocount})$) in the right of Figure 2. This further highlights the dependence of the log-contrast coefficients on the used pseudocount. The fact that the perturbation-based effects explicitly account for zeros and allow us to distinguish between continuous effects (e.g., $\mathrm{CFI}_{\mathrm{unit}}$ or $\mathrm{CFI}_{\mathrm{mult}}$) and presence/absence effects $\mathrm{CKE}$ are an important feature of the perturbation-based approach. In Section S5.8 we provide additional plots which compare the significant hits of $\mathrm{CFI}_{\mathrm{mult}}$, log-contrast estimates and a naive marginal analysis and illustrate the differences between $\mathrm{CFI}_{\mathrm{unit}}$ and $\mathrm{CKE}$ when compared to estimates using $\mathrm{CFI}_{\mathrm{mult}}$.
@@ Acknowledgements
Table S1: Numerical illustration of Examples S1 and S2. The first row in each table corresponds to the marginal treatment effect (MTE or OLS of $Y$ on $-G(Z)$) estimate based on a correctly specified linear model. The second row corresponds to estimates of two of the target parameters given in Table 1 with random forest (rf) as nuisance function estimators. Bold red numbers indicate a significant negative effect at a $5\%$ level, while bold green indicates a significant positive effect. Left: The estimated MTE is negative, even though the true effect is positive. The compositional knock-out effect ($\mathrm{CKE}^{1}$), which adjust for the simplex structure, results in the correct sign. Right: The estimated marginal effect of an increase in diversity (measured by the negative Gini coefficient, i.e., $-G(Z)$) is negative even though the true effect is equal to $1$ by construction. The compositional diversity effect $\mathrm{CDI}_{\mathrm{Gini}}$ correctly recovers the effect. Details on both experiments are provided in Section S5.1.
@@
In this section we provide a general method for finding derivative-isolating reparametrizations based on the development in Section 3.2. Our main result is as follows.
@@
Part (a) starts from a given speed and finds a corresponding derivative-isolating reparametrization by solving the differential equation in (S3). Part (b) starts from a one-dimensional summary statistic, which is assumed to be smoothly decreasing in the distance from $z$ to the endpoint. Using the statistic as one component in a derivative-isolating reparametrization, the result provides the speed of the perturbation that is being parametrized. In the following example, we give the missing details from Examples 1 and 2.
@@
S2.6 Amalgamation effects
In this section we consider general amalgamating perturbations as defined in Section 3.1 rather than only those which push towards a vertex of the simplex, i.e., those we describe in Section 3.3. We first generalize Example 1.
@@
With this generalization, we can consider perturbation effects for general amalgmating perturbations. As an example, consider a study where data is collected on how individuals spend their time each day (e.g., hours spent sleeping, eating, working, doing sports, etc.) as well as a response variable (e.g., general happiness). Assume now that we want to use this data to investigate the relationship between time spent on sports and happiness. As in Section 3.3, we are again interested in the effect of an individual component, so we might consider estimating a $\mathrm{CFI}^{j}$. However, this would capture the effect of increasing time spent on sports while all other activities are down-scaled proportionally. For example, time spent sleeping and eating would both be decreased if time doing sports is increased which is not realistic. To address this, we can instead consider a general $(A\to B)$-amalgamating perturbation (as defined in Section 3.1), where $B=\{\text{sport}\}$ and $A$ corresponds to activities which are reduced to make increased time for sports. Perturbing in this fashion keeps the components in $[d]\setminus(A\cup B)$ fixed (e.g., eating and drinking) while down-scaling the activities in $A$. We call directional perturbation effects of this form compositional amalgamation influences, or $\mathrm{CAI}^{A\to B}$ for short. As with the $\mathrm{CFI}^{j}$, we use subscripts to denote the speed of the perturbation, so we have $\mathrm{CAI}_{\mathrm{unit}}^{A\to B}$ and $\mathrm{CAI}^{A\to B}_{\mathrm{mult}}$ (see Example S5).
Analogously, we could be interested in the effect of removing particular activities (e.g. gambling) entirely and spending the time on desirable activities (e.g. sports) while fixing time spent on a third group of activities (e.g. eating, sleeping). We can capture this effect using a binary amalgamating perturbation $\psi$. We call the resulting average binary perturbation effect for $\psi$ an $(A\to B)$-compositional amalgamation effect, or $\mathrm{CAE}^{A\to B}$ for short.
@@
In this section we describe an alternative way to construct perturbations than the one considered in Sections 3.1 and 3.2. When constructing the directional perturbations in Section 3.2, we considered (Euclidean) straight lines to the endpoint to construct directions, that is, $v_{E}:z\mapsto(E(z)-z)/\|E(z)-z\|_{1}$. While we believe that this is natural in many applications, one could also imagine starting from alternative geometries on the simplex and construct other types of directions. We illustrate how this would work for the Aitchison geometry and show that it provides an interesting connection to the log-ratio based literature. The Aitchison geometry is a popular alternative geometry that goes back to the work of Aitchison [1982] who proposed to analyze the interior of the simplex by transforming to a subset of Euclidean space and applying conventional methods to the transformed data. A popular variant of this transformation is the centered log-ratio transform (CLR) defined for all $z\in(\Delta^{d-1})^{\circ}$ (the interior of the simplex) by
@@
As a first example, assume that the SCM given in Figure S2 (left) generates the data. Given a perturbation $\psi$ that we are interested in, a derivative-isolating reparametrization (see Section 3.2) $\phi:\Delta^{d-1}\rightarrow\mathbb{R}\times\mathcal{W}$ maps $Z$ to $(L,W)$ such that if we intervene on $Z$ by setting it to $\psi(Z,\gamma)$ this corresponds to an intervention in an SCM over the variables $(Y,L,W)$ on $L$ alone (see Figure S2 (right)).
@@
In this section we provide details on the theory and algorithms for the estimation of perturbation effects. Section S4.1 introduces the nonparametric estimators of perturbation effects. Section S4.2 contains a summary of the theoretical results and algorithms for the main estimation procedures. Section S4.3 contains a proof of our main theoretical result. Section S4.4 investigates the robustness of the different estimators. Section S4.5 describes how to modify the estimation of $\tau_{\psi}$ using a derivative-isolating reparametrization in the presence of observations with $\omega_{\psi}(Z)=0$. Section S4.6 describes our procedure for the estimation of the derivative of a generic regression function. Section S4.7 describes our procedure for the estimation of the conditional score function $\rho$ that is required for the estimation of an average partial effect.
S4.1 Nonparametric estimation of perturbation effects
S4.1.1 Binary perturbation effect
The denominator and $\mathbb{E}[Y]$ are both easy to estimate efficiently using their empirical counterparts (the one-step correction term is 00 for these quantities). The term $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ in the numerator is trickier and is related to the notion of the average predictive effect [e.g., Robins et al., 1994, Chernozhukov et al., 2018, Kennedy, 2023] that has been analyzed extensively in the causal inference literature. The one-step corrected estimator for $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ corresponds to the augmented inverse propensity weighted (AIPW) estimator and is given by
$\widehat{\lambda}_{\psi}:=\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{f}(1,W_{i})-Y_{i}+\frac{Y_{i}-\widehat{f}(L_{i},W_{i})}{\widehat{\pi}(W_{i})}L_{i}}{\frac{1}{n}\sum_{i=1}^{n}(1-L_{i})}.$ (S8)
S4.1.2 Directional perturbation effect
This quantity is another well-known target of inference; the average partial effect [Newey and Stoker, 1993, Chernozhukov et al., 2022, Klyne and Shah, 2023]. The one-step corrected estimator in this setting is given by
$\widehat{\tau}_{\psi}:=\frac{1}{n}\sum_{i=1}^{n}\partial_{\ell}\widehat{f}(L_{i},W_{i})-\widehat{\rho}(L_{i},W_{i})(Y_{i}-\widehat{f}(L_{i},W_{i}))$ (S9)
For the estimation of $f$ we require that the estimator $\widehat{f}$ is differentiable and with computable derivatives (which additionally need to approximate the true derivative at a sufficiently fast rate). This excludes the direct use of many popular machine learning methods such as random forests or boosted trees. One solution is to start from a potentially non-differentiable pilot estimate of $f$ and smooth it to get a differentiable estimate for which derivatives can be computed. Such a procedure based on kernel-smoothing has been proposed by Klyne and Shah [2023]. Here, we use an alternative smoothing method which uses random forests and local polynomials. Finally, for the estimation of $\rho$, we follow the proposal in Klyne and Shah [2023] and assume a location-scale model. This requires the estimation of the conditional mean and variance of $L$ given $W$ (referred to as $\eta$ in Algorithm S1). Further details on our smoothing method and the location-scale based estimate of $\rho$ are given in Sections S4.6 and S4.7, respectively.
S4.2 Theoretical analysis of estimators
In this section we present a result on asymptotic normality uniformly over large classes of distributions for each of the three estimation methods described in Section 4; nonparametric estimation of $\tau_{\psi}$, nonparametric estimation of $\lambda_{\psi}$ and by means of estimating $\theta$ in the partially linear model. These results are derived using standard proof techniques from the semiparametric literature and likely exist individually in similar form elsewhere [e.g., Klyne and Shah, 2023, Chernozhukov et al., 2018]. We do not claim that any of the derived asymptotic results are novel or surprising but include them for completeness and to motivate the precise form of the estimators and their corresponding variance estimates.